If the volume of cube is 700 inches^3 and S= length of a side, what is the length of a side? Equation: V= S^3
Answer by oberobic(2304)

S = length of a side
V = volume = S^3 = 700 cubic inches
.
Using a calculator, you can use the exponent of 1/3 to get the cube root...
.
7001/3 = 8.879040017
.
That makes sense because 9^3 = 9*9*9 = 729.
In contrast, 8^3 = 8*8*8 = 512.
So, logically the cube root should be much closer to 9 than to 8.
.
The length of each side is 8.879040017.
.
Done